# Voltage Support Procurement in Transmission Grids: Incentive Design via Online Bilevel Games

### v1

@@ I Introduction
In this paper, we first review the motivating case study of the Swiss voltage procurement scheme (Section II). Then, in Section III, we use game theory’s formalism to provide a rigorous mathematical formulation of the problem faced by the TSO (in deciding the incentives) and the DSOs (in responding optimally). In Section IV we presente the main result: we consider the case in which incentives can be updated in real time and propose a computational solution that allows the TSO to adjust incentives based on the DSO response so that the desired grid voltage profile can be achieved. The key tool in this design is adapted from [15] and consists of a gradient-based iterative algorithm for the optimal intervention in Stackelberg games. We finally illustrate the performance of this hierarchical service procurement solution in simulations (Section V) and discuss the remaining open challenges in the Conclusions.
@@ II A motivating case study: The Swiss voltage support procurement scheme
The analysis of historical data provided by Swissgrid (Figure 5) indicates that reactive power responses from DSOs do not consistently result in positive financial rewards and often lead to significant voltage-tracking errors. An exception is represented by traditional power plants, which typically achieve near-perfect voltage tracking, except when limited by their reactive power constraints. It also appears from the data that many DSOs may frequently reach their reactive power limits, resulting in penalties. However, these constraints are currently not incorporated into the OPF-V program, determining reference profiles $v_{\text{ref}}$ that are infeasible to track.
@@ III Game-theoretic formulation
where $R,X$ are derived from the resistance and reactance matrices, and $v_{0}$ is the baseline voltage profile of the system in the absence of any power consumptions. A standard assumption is that the diagonal elements of $X$ (the equivalent grid reactance seen by each bus) are negative.
@@ III-A Properties of the lower level game (DSOs)
where $X_{ii}$ represents the self-sensitivity of voltage at bus $i$ to its own reactive power consumption, and $X_{ij}$ is the mutual-sensitivity. Substituting it into the objective function, we have:
$f_{i}(v_{i}^{\text{ref}},\xi_{i},\xi_{-i})=\frac{1}{2}\left(C_{i}-\gamma X_{ii}\right)\xi_{i}^{2}-\gamma K\xi_{i},$ (4)
@@
Assume $\lambda_{\min}(C-\gamma X)\geq\mu>0$. Then, the pseudo-gradient mapping $F(v_{\mathrm{ref}},\cdot)$ is $\mu$-strongly monotone and $L_{F}$-Lipschitz continuous, with $L_{F}=\|C-\gamma X\|$.
@@
$F(v_{\text{ref}},\xi)=(C-\gamma X)\xi-\gamma(v_{0}+Rp-v_{\text{ref}}).$
where $C$ is the diagonal matrix of cost coefficients.
@@
$J_{F}\coloneqq\nabla_{\xi}F(v_{\text{ref}},\xi)=C-\gamma X.$
Because both $C$ and $X$ are symmetric, $J_{F}$ is a symmetric matrix.
Since $\lambda_{\min}(C-\gamma X)\geq\mu>0$, the Jacobian $J_{F}$ is positive definite. Consequently, for any $\xi_{1},\xi_{2}$, we have:
@@
$\|F(v_{\text{ref}},\xi_{1})-F(v_{\text{ref}},\xi_{2})\|\leq\|C-\gamma X\|\|\xi_{1}-\xi_{2}\|,\quad\forall\xi_{1},\xi_{2}.$
Denoting the upper bound of the constant matrix norm by $L_{F}\coloneqq\|C-\gamma X\|$, we conclude that $F(v_{\text{ref}},\cdot)$ is $L_{F}$-Lipschitz continuous. ∎
@@
The condition $\lambda_{\min}(C-\gamma X)\geq\mu>0$ can be enforces in the design of the incentive by selecting $\gamma$ sufficiently small. It is also automatically verified if $X$ is negative definite, which is the case for example of radial networks [16, Lemma 1].
@@ IV Co-design of incentives and DSOs’ response
Based on this game-theoretic formulation, we propose an online protocol to update $v_{\text{ref}}$ as a dynamic incentive signal, and simultaneously automate the response of the DSOs, so that they jointly converge to the optimal grid operation defined by the Stackleberg game in (1).
@@ IV-D Algorithm convergence
Finally, we note that a constant step size $\epsilon$ would be required in (19) to ensure continuous incentive updates in online settings, where the parameters of the bilevel game (1) (e.g., power injections $p$) vary over time. However, the responsiveness of the online implementation to time-varying conditions benefits from non-diminishing step sizes. The development of a formal stability and tracking analysis for this case is left for future work.

### v2

@@ I Introduction
In this paper, we first review the motivating case study of the Swiss voltage procurement scheme (Section II). Then, in Section III, we use game theory’s formalism to provide a rigorous mathematical formulation of the problem faced by the TSO (in deciding the incentives) and the DSOs (in responding optimally). In Section IV we present the main result: we consider the case in which incentives can be updated in real time and propose a computational solution that allows the TSO to adjust incentives based on the DSO response so that the desired grid voltage profile can be achieved. The key tool in this design is adapted from [15] and consists of a gradient-based iterative algorithm for the optimal intervention in Stackelberg games. We finally illustrate the performance of this hierarchical service procurement solution in simulations (Section V) and discuss the remaining open challenges in the Conclusions.
@@ II A motivating case study: The Swiss voltage support procurement scheme
The analysis of historical data provided by Swissgrid (Figure 5) indicates that reactive power responses from DSOs do not consistently result in positive financial rewards and often lead to significant voltage-tracking errors. An exception is represented by traditional power plants, which typically achieve near-perfect voltage tracking, except when limited by their reactive power constraints. It also appears from the data that many DSOs may frequently reach their reactive power limits, resulting in penalties. However, these constraints are currently not incorporated into the OPFV program, determining reference profiles $v_{\text{ref}}$ that are infeasible to track.
@@ III Game-theoretic formulation
where $R,X$ are derived from the resistance and reactance matrices, and $v_{0}$ is the baseline voltage profile of the system in the absence of any power consumptions.
@@ III-A Properties of the lower level game (DSOs)
where $X_{ii}$ represents the self-sensitivity of voltage at bus $i$ to its own reactive power consumption, and $X_{ij}$ is the mutual-sensitivity. Substituting $v_{i}$ into the objective function, we have:
$f_{i}(v_{i}^{\text{ref}},\xi_{i},\xi_{-i})=\frac{1}{2}\left(C_{i}-2\gamma X_{ii}\right)\xi_{i}^{2}-\gamma K\xi_{i},$ (4)
@@
Assume $\lambda_{\min}(C-\gamma\tilde{X})\geq\mu>0$. Then, the pseudo-gradient mapping $F(v_{\mathrm{ref}},\cdot)$ is $\mu$-strongly monotone and $L_{F}$-Lipschitz continuous, with $L_{F}=\|C-\gamma\tilde{X}\|$.
@@
$F(v_{\text{ref}},\xi)=(C-\gamma\tilde{X})\xi-\gamma(v_{0}+Rp-v_{\text{ref}}),$
where $C$ is the diagonal matrix of cost coefficients and $\tilde{X}\coloneqq X+\operatorname{diag}(X_{ii})_{i\in\mathcal{N}}$.
@@
$J_{F}\coloneqq\nabla_{\xi}F(v_{\text{ref}},\xi)=C-\gamma\tilde{X}$.
Because both $C$ and $\tilde{X}$ are symmetric, $J_{F}$ is a symmetric matrix.
Since $\lambda_{\min}(C-\gamma\tilde{X})\geq\mu>0$, the Jacobian $J_{F}$ is positive definite. Consequently, for any $\xi_{1},\xi_{2}$, we have:
@@
$\|F(v_{\text{ref}},\xi_{1})-F(v_{\text{ref}},\xi_{2})\|\leq\|C-\gamma\tilde{X}\|\|\xi_{1}-\xi_{2}\|,\quad\forall\xi_{1},\xi_{2}.$
Denoting the upper bound of the constant matrix norm by $L_{F}\coloneqq\|C-\gamma\tilde{X}\|$, we conclude that $F(v_{\text{ref}},\cdot)$ is $L_{F}$-Lipschitz continuous. ∎
@@
The condition $\lambda_{\min}(C-\gamma\tilde{X})\geq\mu>0$ is automatically verified if $\tilde{X}$ is negative definite, which is the case of radial networks [16, Lemma 1]. For meshed network where $\tilde{X}$ is not negative definite, the condition can be enforced in the design of the incentive by selecting $\gamma<\frac{c_{\min}}{\lambda_{\max}\left(\tilde{X}\right)}$.
@@ IV Co-design of incentives and DSOs’ response
Based on this game-theoretic formulation, we propose an online protocol to update $v_{\text{ref}}$ as a dynamic incentive signal, and simultaneously automate the response of the DSOs, so that they jointly converge to the optimal grid operation defined by the Stackelberg game in (1).
@@ IV-D Algorithm convergence
Finally, we note that a constant step size $\epsilon$ would be required in (19) to ensure continuous incentive updates in online settings, where the parameters of the bilevel game (1) (e.g., power injections $p$) vary over time. The development of a formal stability and tracking analysis for this case is left for future work.